# From Large to Super-Tiny: End-to-End Optimization for Cost-Efficient LLMs

### v2

@@ Abstract
In recent years, Large Language Models (LLMs) have significantly advanced artificial intelligence by optimizing traditional Natural Language Processing (NLP) pipelines, improving performance and generalization. This has spurred their integration into various systems. Many NLP systems, including ours, employ a "one-stage" pipeline directly incorporating LLMs. While effective, this approach incurs substantial costs and latency due to the need for large model parameters to achieve satisfactory outcomes. This paper introduces a three-stage cost-efficient end-to-end LLM deployment pipeline—including prototyping, knowledge transfer, and model compression-to tackle the cost-performance dilemma in LLM-based frameworks. Our approach yields a super tiny model optimized for cost and performance in online systems, simplifying the system architecture. Initially, by transforming complex tasks into a function call-based LLM-driven pipeline, an optimal performance prototype system is constructed to produce high-quality data as a teacher model. The second stage combines techniques like rejection fine-tuning, reinforcement learning, and knowledge distillation to transfer knowledge to a smaller 0.5B student model, delivering effective performance at minimal cost. The final stage applies quantization and pruning to extremely compress models to 0.4B, achieving ultra-low latency and cost. The framework’s modular design and cross-domain capabilities suggest potential applicability in other NLP areas.
* Equal Contribution Corresponding Author
@@ 1 Introduction
Figure 1: From the traditional BERT-based system to the LLM-based system.
@@
In recent years, Large Language Models (LLMs) have made significant strides [1; 2; 3] in enhancing artificial intelligence by addressing shortcomings of traditional pipelines — including manual annotation, suboptimal performance, and limited generalization and adaptability. Consequently, LLM-based pipelines have stimulated substantial demand for integration into various Natural Language Processing (NLP) systems. We find that numerous NLP systems, including our own, have directly adopted LLM into the pipeline, often referred to as the "one stage" pipeline. Despite the breakthroughs facilitated by this pipeline, it also incurs substantial cost increases [4] and high inference latency [5]. Within our NLP system, cost-performance is an especially critical metric, a sentiment likely shared by others in the field [6]. These significant costs limit the broader applicability of LLMs and impact their commercial viability. Our analysis points out that it is precisely the simplistic training method of the one-stage pipeline, which necessitates larger model parameters to achieve satisfactory results, that leads to significant costs.
@@
In this paper, we introduce a three-stage performant and cost-efficient end-to-end LLM deployment pipeline, consisting of prototyping, knowledge transfer, and model compression, to address the pervasive "cost-performance" dilemma in the practical deployment of LLM-based frameworks. This approach ultimately yields a super tiny model, which is deployed in our online system, achieving optimal cost-performance metrics while substantially simplifying system architecture. The first stage involves constructing the prototype system, where traditional complex tasks are converted into function call-based LLM-driven workflows [7] through prompt engineering, resulting in an optimal performance prototype compared to traditional pipeline. This prototype serves as a teacher model in the second stage, efficiently generating vast quantities of high-quality data. During the knowledge transfer phase, leveraging the optimal prototype from the first stage, we employ techniques including rejection fine-tuning (RFT), reinforcement learning (RL) [8; 9; 10], and multi-strategy knowledge distillation (KD) targeting the Qwen 2.5 series model [11], effectively transferring knowledge to a 0.5B student model. The derived student model is considerably smaller than the prototype model while achieving nearly comparable performance at minimal cost. In the third stage, we utilize quantization and pruning techniques to further compress the student model developed in the second phase, resulting in a remarkably compact model of 0.4 billion parameters (as shown in Figure 1). Ultimately, this super tiny model is deployed in our online system. This model achieves hundreds-fold compression and extremely low latency and cost in exchange for a trivial and acceptable loss in performance. The modular design and superior cross-domain performance of the framework suggest its potential applicability in other NLP domains. The major contributions of this paper are as follows:
@@
  • The first validated three-stage cost-efficient end-to-end LLM pipeline with simplifying original system complexity achieves SOTA results and resolves the "cost-performance" dilemma, realizing a 180-fold compression with nearly consistent AR compared to 72B LLMs and up to an absolute 14% AR increase compared to traditional BERT-based pipelines.
@@
  • We propose the first hybrid strategy that significantly enhances the generalization of super tiny models on domain-specific tasks through combining RL and KD. This approach effectively transfers the capabilities of LLMs to smaller models.
@@ 2 Related Works
Leveraging in-context learning and reasoning capabilities, LLMs can select functions and parameters from provided lists based on user intent [12]. Fine-tuned 7B models like Gorilla [13] achieve GPT-4-level Function Call accuracy on datasets with thousands of APIs. Building on this, TinyAgent [14] enables function calling capabilities for edge devices with sub-1.1B parameter models. Query-to-API reformulation via rejection sampling can also optimize API retrieval, reducing retrieval latency while improving accuracy [15]. Nowadays, LLMs can be used to evaluate model-generated content [16; 17]. CriticGPT [18] indicates that LLMs can build a cost-effective and high-precision judge, surpassing humans in detecting subtle errors.
@@
Knowledge distillation (KD) [19] transfers knowledge from a teacher model to a smaller student model by simulating probability distribution. Common approaches include Logits-based[20; 21], Feature-based [22] and sequence-level (Seq-KD) [23] distillation. Initially, Logits-based KD uses the forward KL divergence (FKL) [24; 25] as the loss function. Some KD methods prefer Reverse KL divergence (RKL) [26; 27; 28; 21]. Recently, Adaptive Kullback-Leiber (AKL) divergence [29] achieves a balance between mimicking distribution and avoiding overfitting.
@@
Efficient LLM deployment primarily relies on pruning[30; 31] and quantization[32; 33; 34]. Structured pruning[35] employs layer removal[36; 37] or dimensionality reduction[31; 37; 38] using gradient-based[31; 38] or activation-aware[37] importance criteria, implemented via learnable masks[38; 39] or pruning-retraining pipelines[36; 31; 37]. Weight-only LLM quantization methods[32; 33; 40] improve memory efficiency while joint weight-activation schemes[34; 41] and their combination [42] accelerate computation.
@@
Recent works [8; 9; 10] focus on leveraging Reinforcement Learning (RL) to enhance model reasoning capabilities. Notable reasoning models such as DeepSeek-R1 [43], QWQ-32B [44], and OpenAI o1 [45] have emerged, employing RL algorithms like Proximal Policy Optimization (PPO) [46], Group Relative Policy Optimization (GRPO) [47], and others [48; 49; 50]. For general-purpose LLMs, preference optimization algorithms [51; 52; 53; 54] are widely adopted. Additionally, self-improving pipelines [55; 56; 57] have been proposed to iteratively refine model performance. We primarily utilize RL to enhance the model’s generalization. Compared to directly distilling a reasoning model, RL results in shorter and more precise Chains-of-Thought (CoT).
@@ 3 Methodology
Figure 2: Overview of Prototype, Production and Application.
Recent advancements in LLMs have facilitated significant progress in NLP tasks such as classification and summarization. We harness these capabilities to reconstruct the traditional dialogue system through a cost-efficient end-to-end LLM pipeline, which includes prototyping, knowledge transfer, and model compression, as illustrated in 2(a) and 2(b). Initially, we develop the prototype system by leveraging the exceptional performance of contemporary LLMs. Subsequently, we perform knowledge transfer from large models to smaller models. Finally, we achieve additional reduction in model size through further model compression techniques. Ultimately, our objective is to deliver intelligent user experiences at minimal costs, as demonstrated in 2(c).
@@ 3.1 Prototyping
Table 1: Complexity statistics of Our dataset vs Open Source datasets. More details in Appendix A.
Dataset Test Set Size Intents Slots TTR Entropy
TinyAgent[14] 1023 17 48 0.0087 7.8220
Snips[58] 893 7 72 0.0207 7.6997
ATIS[59] 700 21 120 0.0111 7.4148
Our dataset 11922 74 225 0.1410 9.4250
@@
Traditional dialogue systems usually employ the Intent Classification and Slot Filling (ICSF) framework, often utilizing BERT[60; 61]. However, a single model often struggles to handle complex real-world issues, as shown in the Table 1. Consequently, a Domain Classifier (DC) is frequently used to first categorize queries into domains, enabling domain-specific ICSF parsing. Although the hierarchical DC-ICSF architecture improves accuracy, the system complexity increases with feature updates. This leads to high inter-module coupling, increasing development and maintenance costs. Furthermore, ambiguous domain boundaries turn the DC into a bottleneck for overall end-to-end accuracy, creating a "feature expansion - operational difficulty" feedback loop that hinders dialogue experience optimization.
Modern LLMs exhibit adequate capability to tackle these NLP problems. Particularly in related Tool Use works, LLMs can learn to call functions without additional training by attaching function descriptions in prompts. Essentially, ICSF can be viewed as a function call task, where the model extracts critical information from user queries as function parameters for specific modules to execute. Following this approach, we design a prototype system based on Qwen2.5-72B-Instruct[11]. As illustrated in 2(a), the prototype system accepts user queries and contextual information (including chat history and temporal status information) as input, searches relevant APIs from the API pool, then outputs function names and parameters to call. The traditional system comprises one DC module and multiple ICSF models (based on BERT or regular expressions), each requiring separate maintenance and frequent DC updates. In contrast, the LLM-based system, equipped with function call capability, processes all tasks through a unified model, as illustrated in Figure 1, thereby reducing system complexity and enhancing operational efficiency.
To enhance temporal awareness and mitigate hallucinations, we enrich the input with contextual knowledge via RAG [62], retrieving explanations and extensions of entity-related terms from a curated knowledge base. Instead of adding all callable APIs from the API pool into the LLM prompt, we implement an API retrieval module inspired by [15]. The LLM first generates potential API descriptions, then retrieves top-K APIs via vector similarity (using GTE-Qwen2-1.5B [63] as the embedding model and FAISS [64] for indexing). Then, the LLM selects optimal functions/parameters from candidate APIs. If no suitable function exists, the system returns "No appropriate function"; in the case of ambiguous queries, "Request unclear" are generated. For backward compatibility, we design an Adapter to map functions/parameters to legacy domain-intent-slot structures.
@@
The traditional evaluation scheme for ICSF relies on the test set that requires full alignment of Intent and Slots. However, in real-world scenarios, there exist multiple solutions for function and slot selection. For example: (1) "Slightly increase the volume" could be resolved using either volume setting or volume increasing APIs; (2) "11:00 pm" and "23:00" semantically represent the same time. Additionally, due to crowdsourced annotators’ limited understanding of specific business logic, the manual labeling accuracy remains suboptimal. As shown in 2(c), the crowdsourced labeler’s annotation accuracy has become the system bottleneck.
To address issues above, we introduce LLM-as-a-judge to evaluate prototype accuracy. We design a Judge based on QwQ-32B[44]: Given context and a set of Function & parameters, the Judge determines whether this combination semantically satisfies the user query. Following Samuylova’s best practices [^1] https://www.evidentlyai.com/llm-guide/llm-as-a-judge, we optimize the Judge using a 1k-expert-annotated dataset. Subsequently, both the LLM and 9 domain experts evaluate 3k test samples under identical conditions. Results show that Judge can provide logically coherent Chain-of-Thought(CoT) rationales and it achieves 93% consistency with expert judgments. Based on these findings, we conclude that LLM-as-a-Judge is the optimal evaluation approach for this task. As demonstrated in Figure 4, we reconstruct the dialogue pipeline from BERT-based DC-ICSF architecture to LLM-based Function Call architecture. Evaluation through LLM-as-a-Judge demonstrates that the new architecture improves accuracy from 81.43% to 94.74% compared to the original one.
@@ 3.2 Knowledge Transfer
Building on our agent-oriented large model design, we propose Supervised Fine-Tuning (SFT) on smaller models using large model-generated data instead of relying solely on in-context learning (ICL) with large models. This approach offers two key benefits: (1) Smaller models can achieve performance comparable to large models after trained on sufficient high-quality data; (2) It significantly boosts inference and optimization efficiency. For a dataset $D$ containing prompt component $x$ and response component $y$, we define $D(x,y)=\{x_{i},y_{i}\}_{i=1}^{N}$, where $N$ denotes the dataset size. To minimize the Cross Entropy (CE) loss, we first utilize the SFT optimization, and then employ further optimization Rejection sampling Fine-Tuning (RFT). We use LLM judge introduced in Section 3.1 to evaluate training data generated by the LLM prototype, filter out low-quality data, and generate the final training set. The actual training data is formulated as:
@@
Reinforcement Learning
Since OpenAI introduced Reinforcement Learning from Human Feedback (RLHF) [65], it has become an essential component for aligning LLMs, as evidenced in major technical reports [66; 67; 68; 11; 69]. DeepSeek-R1 [43] proposes a cost-effective RL training method from scratch which combined rule-based ORM and GRPO [47]. Subsequent works like Logic-RL [8; 9; 10] applies this efficient approach across various tasks, enabling the model to automatically generate Chain of Thought (CoT) during training, thereby enhancing performance [70]. Based on our observation that model generalization degrades after RFT, we attempt to employ RL to improve model performance on out-of-domain datasets. Following Logic-RL [9], we refine the reward design for our specific task (see Appendix B) and conduct extensive attempts to integrate knowledge distillation (see Appendix K).
Distillation
Using the Function Call-related data from large models and the RFT methods, we transfer capabilities from large to small models through training. To accelerate model training and inference, we implement prompt compression during training by removing unnecessary content from input-output pairs, including format specifications in inputs and Chain-of-Thought (CoT) rationales in outputs. Please refer to Appendix C for details.
Knowledge distillation [19] is an effective method for transferring knowledge from large to small models, requiring fewer data and exhibiting lower overfitting risks compared to SFT. We adopt the logits-based Adaptive Kullback-Leibler (AKL) distillation [29], with the optimization objective:

### v3

@@ Abstract
Large Language Models (LLMs) have significantly advanced artificial intelligence by optimizing traditional Natural Language Processing (NLP) workflows, facilitating their integration into various systems. Many such NLP systems, including ours, directly incorporate LLMs. However, this approach either results in expensive costs or yields suboptimal performance after fine-tuning. In this paper, we introduce a three-stage cost-efficient end-to-end LLM deployment pipeline, comprising prototyping, knowledge transfer, and model compression, to effectively tackle the cost-performance dilemma in LLM-based frameworks. Its high cost-efficiency is manifested not only in simplifying system complexity and producing super-tiny online models with enhanced performance and reduced costs in the results, but also in addressing development cycle constraints, the lack of extensive high-quality data, and limited computational resources during the project development process. In the first stage, we construct an optimal performance prototype system by transforming complex tasks into a function call-based LLM-driven pipeline, which serves as a teacher model to generate high-quality data. In the second stage, we combine techniques like rejection sampling fine-tuning, reinforcement learning, and knowledge distillation to transfer knowledge to 0.5B student models, delivering effective performance at minimal cost. In the final stage, we further compress models to 0.4B via quantization and pruning, achieving ultra-low latency and cost. Extensive experimental results and the framework’s modular design suggest cross-domain capabilities and potential applicability in other NLP areas.
[…]
@@ 1 Introduction
In recent years, Large Language Models (LLMs) have made significant strides [1, 2, 3] in enhancing artificial intelligence by addressing shortcomings of traditional pipelines — including manual annotation, suboptimal performance, and limited generalization and adaptability. Consequently, LLM-based pipelines have stimulated substantial demand for integration into various Natural Language Processing (NLP) systems. We find that numerous such NLP systems, including our own, have directly utilized APIs of closed-source LLMs or fine-tuned open-source LLMs, which can be referred to as the "one-stage" pipeline. Although this pipeline offers implementation simplicity, it suffers from substantial costs [4] and high inference latency [5] or suboptimal performance, which limit the broader applicability of LLMs and impact their commercial viability. Within our NLP system, cost-performance is an especially critical metric, a sentiment likely shared by others in the field [6]. Moreover, our analysis indicates that this pipeline struggles to balance the cost and the performance due to insufficient high-quality data and monotonous optimization techniques.
In this paper, we introduce a three-stage cost-efficient end-to-end LLM deployment pipeline, as shown in Figure 1, consisting of prototyping, knowledge transfer, and model compression, to address the pervasive "cost-performance" dilemma in the practical deployment of LLM-based frameworks. The high cost-efficiency of this approach is demonstrated by both the outcomes and the process. In terms of outcomes, the approach simplifies system complexity and produces super-tiny models with optimal performance and reduced inference costs and latency. Regarding the process, the approach addresses challenges such as constrained development cycles, the scarcity of substantial high-quality data, and limited computational resources.
@@
In our proposed pipeline, the first stage involves constructing the prototype system, where traditional complex tasks are converted into function call-based LLM-driven workflows [7] through prompt engineering, resulting in an optimal performance prototype compared to traditional pipeline. This prototype serves as a teacher model in the second stage, efficiently generating vast quantities of high-quality data. In the second stage, leveraging the optimal prototype from the first stage, we employ techniques including Rejection sampling Fine-Tuning (RFT), Reinforcement Learning (RL) [8, 9, 10], and multi-strategy knowledge distillation (KD) on the Qwen 2.5 series model [11]. Our hybrid strategy effectively transfers generalization capabilities and the abilities of generating high-quality CoT to a 0.5B student model. The derived student model is considerably smaller than the prototype model while achieving nearly comparable performance at minimal cost. In the third stage, we utilize quantization and pruning techniques to further compress the student model developed in the second phase, resulting in a remarkably compact 0.4B model. Ultimately, this super-tiny model is deployed in our online system. This model achieves hundreds-fold compression and extremely low latency and cost in exchange for a trivial and acceptable loss in performance. The modular design and superior cross-domain performance of the framework suggest its potential applicability in other NLP domains. Our major contributions are as follows:
@@
  • The first validated three-stage cost-efficient end-to-end LLM pipeline, simplifying system complexity, achieves SOTA results and resolves the "cost-performance" dilemma. It realizes 180 times compression with nearly consistent performance compared to original LLMs and up to an absolute 14% performance increase over traditional BERT-based systems.
@@
  • We propose the first hybrid knowledge transfer method by combining RL and KD. It effectively transfers the capabilities of LLMs to super-tiny models and achieves performance comparable and even superior to certain larger models on domain-specific tasks.
@@
Figure 1: Workflows of prototyping, knowledge transfer and compression.
@@ 2 Related Works
Leveraging in-context learning and reasoning capabilities, LLMs can select functions and parameters from provided lists based on user intent [12]. Fine-tuned 7B models like Gorilla [13] achieve GPT-4-level Function Call accuracy on datasets with thousands of APIs. Building on this, TinyAgent [14] enables function calling capabilities for edge devices with sub-1.1B parameter models. Query-to-API reformulation via rejection sampling can also optimize API retrieval, reducing retrieval latency while improving accuracy [15]. Nowadays, LLMs can be used to evaluate model-generated content [16, 17]. CriticGPT [18] indicates that LLMs can build a cost-effective and high-precision judge, surpassing humans in detecting subtle errors.
@@
Knowledge distillation (KD) [19] transfers knowledge from a teacher model to a smaller student model by simulating probability distribution. Common approaches include Logits-based[20, 21], Feature-based [22] and sequence-level (Seq-KD) [23] distillation. Initially, Logits-based KD uses the forward KL divergence (FKL) [24, 25] as the loss function. Some KD methods prefer Reverse KL divergence (RKL) [26, 27, 28, 21]. Recently, Adaptive Kullback-Leiber (AKL) divergence [29] achieves a balance between mimicking distribution and avoiding overfitting.
@@
Efficient LLM deployment primarily relies on pruning[30, 31] and quantization[32, 33, 34]. Structured pruning[35] employs layer removal[36, 37] or dimensionality reduction[31, 37, 38] using gradient-based[31, 38] or activation-aware[37] importance criteria, implemented via learnable masks[38, 39] or pruning-retraining pipelines[36, 31, 37]. Weight-only LLM quantization methods[32, 33, 40] improve memory efficiency while joint weight-activation schemes[34, 41] and their combination [42] accelerate computation.
@@
Recent works [8, 9, 10] focus on leveraging Reinforcement Learning (RL) to enhance model reasoning capabilities. Notable reasoning models such as DeepSeek-R1 [43], QWQ-32B [44], and OpenAI o1 [45] have emerged, employing RL algorithms like Proximal Policy Optimization (PPO) [46], Group Relative Policy Optimization (GRPO) [47], and others [48, 49, 50]. For general-purpose LLMs, preference optimization algorithms [51, 52, 53, 54] are widely adopted. Additionally, self-improving pipelines [55, 56, 57] have been proposed to iteratively refine model performance. We primarily utilize RL to enhance the model’s generalization. Compared to directly distilling a reasoning model, RL results in shorter and more precise Chains-of-Thought (CoT).
@@ 3 Methodology
Recent advancements in LLMs have facilitated significant progress in NLP tasks such as classification and summarization. We harness these capabilities to reconstruct the traditional dialogue system through a cost-efficient end-to-end LLM pipeline, which includes prototyping, knowledge transfer, and model compression, as illustrated in 1(a) and 1(b). Initially, we develop the prototype system by leveraging the exceptional performance of contemporary LLMs. Subsequently, we perform knowledge transfer from large models to smaller models. Finally, we achieve additional reduction in model size through further model compression techniques. Ultimately, our objective is to deliver intelligent user experiences at minimal costs, as demonstrated in 1(c).
@@ 3.1 Prototyping
Traditional dialogue systems usually employ the Intent Classification and Slot Filling (ICSF) framework, often utilizing BERT[58, 59]. Given the limitations of a single model in handling complex real-world issues, a Domain Classifier (DC) is often utilized for initial query categorization to facilitate domain-specific ICSF parsing. Although the hierarchical DC-ICSF architecture improves accuracy, the system complexity increases with feature updates. This leads to high inter-module coupling, increasing development and maintenance costs. Furthermore, ambiguous domain boundaries turn the DC into a bottleneck for overall end-to-end accuracy, creating a "feature expansion - operational difficulty" feedback loop that hinders dialogue experience optimization.
Modern LLMs exhibit adequate capability to tackle these NLP problems. Particularly in related Tool Use works, LLMs can learn to call functions without additional training by attaching function descriptions in prompts. Essentially, ICSF can be viewed as a function call task, where the model extracts critical information from user queries as function parameters for specific modules to execute. Following this approach, we design a prototype system based on Qwen2.5-72B-Instruct[11]. As illustrated in 1(a), the prototype system accepts user queries and contextual information (including chat history and temporal status information) as input, searches relevant APIs from the API pool, then outputs function names and parameters to call. The traditional system comprises one DC module and multiple ICSF models (based on BERT or regular expressions), each requiring separate maintenance and frequent DC updates. In contrast, the LLM-based system, equipped with function call capability, processes all tasks through a unified model, thereby reducing system complexity and enhancing operational efficiency.
To enhance temporal awareness and mitigate hallucinations, we enrich the input with contextual knowledge via RAG [60], retrieving explanations and extensions of entity-related terms from a curated knowledge base. Instead of adding all callable APIs from the API pool into the LLM prompt, we implement an API retrieval module inspired by [15]. The LLM first generates potential API descriptions, then retrieves top-K APIs via vector similarity (using GTE-Qwen2-1.5B [61] as the embedding model and FAISS [62] for indexing). Then, the LLM selects optimal functions/parameters from candidate APIs. For backward compatibility, we map functions/parameters to legacy domain-intent-slot structures.
@@
Traditional ICSF evaluation relies on test sets demanding strict intent-slot alignment, contrasting with real-world variability where utterances permit multiple valid interpretations. For example, "Slightly increase the volume" allows alternative API calls, and "11:00 pm"/"23:00" are semantically equivalent. This inherent ambiguity and complex business logic significantly impede consistent human annotation for accurate ground truth.
To address issues above, we introduce LLM-as-a-Judge to evaluate prototype accuracy. We design a judge based on QwQ-32B[44]: Given context and a set of Function & parameters, the judge determines whether this combination semantically satisfies the user query. Following Samuylova’s best practices [^1] https://www.evidentlyai.com/llm-guide/llm-as-a-judge, we optimize the judge using a 1k-expert-annotated dataset. Subsequently, both the LLM and 9 domain experts evaluate 3k test samples under identical conditions. Results show that judge can provide logically coherent Chain-of-Thought(CoT) rationales and it achieves 93% consistency with expert judgments. Based on these findings, we conclude that LLM-as-a-Judge is the optimal evaluation approach for this task. We reconstruct the dialogue pipeline from BERT-based DC-ICSF architecture to LLM-based Function Call architecture. Evaluation through LLM-as-a-Judge demonstrates that the new architecture improves accuracy from 81.43% to 94.74% compared to the original one.
@@ 3.2 Knowledge Transfer
Building on our agent-oriented large model design, we propose Supervised Fine-Tuning (SFT) on smaller models using large model-generated data instead of relying solely on in-context learning (ICL) with large models. This approach offers two key benefits: (1) Smaller models can achieve performance comparable to large models after trained on sufficient high-quality data; (2) It significantly boosts inference and optimization efficiency. For a dataset $D$ containing prompt component $x$ and response component $y$, we define $D(x,y)=\{x_{i},y_{i}\}_{i=1}^{N}$, where $N$ denotes the dataset size.
SFT employs Cross Entropy (CE) loss [63] as the optimization objective. Building on this, we apply Rejection sampling Fine-Tuning (RFT), utilizing an LLM judge, as introduced in Section 3.1, to assess training data generated by the LLM prototype, filter out low-quality data, and construct the final training dataset. The actual training data is formulated as:
@@
Knowledge Distillation
Knowledge Distillation (KD) [19] offers an alternative to RFT for transferring capabilities from large to small models, typically requiring fewer data samples and exhibiting lower overfitting risks. We adopt the logits-based Adaptive Kullback-Leibler (AKL) distillation [29]:
@@
Hybrid Knowledge Transfer
Figure 2: The workflow of the hybrid knowledge transfer strategy.
Following the standard knowledge transfer strategy, we employ either RFT or KD combined with prompt compression, which removes unnecessary format specifications and Chain-of-Thought (CoT) rationales (see Appendix B), to create a fast but domain-specific model. However, our experiments reveal limitations of this conventional approach. Utilizing RFT or KD alone can improve inference efficiency but fails to enhance out-of-domain generalization. According to Logic-RL [9] with task-specific reward adjustment (see Appendix A), we employ Reinforcement Learning (RL) to improve generalization. Additional experiments indicate that employing RL alone generally leads to varying effectiveness across different model sizes: RL training yields exceptional results on small models (7B), surpassing both the LLM prototype and RFT-trained models. On the other hand, when applied to super-tiny models (0.5B), RL becomes ineffective as these models tend to generate shorter or meaningless CoT responses, leading to degraded output diversity.
These findings motivate us to propose a hybrid knowledge transfer strategy, as illustrated in Figure 2. This hybrid strategy first employs RL to train models that are as small as possible while ensuring robust generalization and the ability to generate high-quality CoT, exemplified by the 7B RL-trained model in our experiments. Subsequently, KD is used to transfer these generalization capabilities and CoT generation abilities to super-tiny models. With CoT generation abilities, super-tiny models can be trained through a second round of RL to further enhance generalization. Experimental results demonstrate that super-tiny models exhibit impressive improvement in both in-domain and out-of-domain performance after the second RL training. Detailed performance comparison of different transfer strategies shows in Appendix J.